# Complex variational autoencoders admit Kähler structure

### v3

@@ 2 Notation and conventions
to denote the complex Hessian with Dolbeault operators, i.e. it is a differential operator on smooth sections. We will denote
@@
are the Wirtinger derivatives and $\mu$ is a mean vector (we are not saying all mixed Hessians vanish). We will show this is empirically supported in Appendix C, Figures 6, 7. In particular, we find our architectures work well with this criterion. We specifically do not restrict our decoder to be holomorphic since it is well known a real-valued function that is holomorphic over a connected and open domain is constant due to the Cauchy-Riemann equations. In practice, we will compute Wirtinger derivatives using real-valued surrogates in our VAEs for simplicity reasons. For complex integration, we will use the fact that complex integration has an equivalent form over $\mathbb{R}^{p}$
@@ 3 The complex variational autoencoder (VAE)
$\psi_{\text{Im}}=\frac{\sigma+\delta}{2\sigma+2\text{Re}(\delta)},\psi_{\text{Re}}=i\frac{\sqrt{\sigma^{2}-|\delta|^{2}}}{2\sigma+2\text{Re}(\delta)}.$ (11)
@@ 4 Euclidean VAEs unveil Riemannian structure
(under classical regularity conditions such as absolute continuity, and slight abuse of notation on the subscripts). Thus the sampling is done with the constructed $G$ (or a valid surrogate or replacement). By sampling according to this distribution, they achieved a way of sampling with correspondence to the model geometry, which has shown empirical success in achieving high-fidelity generation chadebec2022geometricperspectivevariationalautoencoders. The arguments we will present will be in the context of Fisher information statistical theory. Our experiments, established in an alternative view, will not attempt to draw from a similar distribution for FID reasons as they do in this work.
@@
$\partial_{z_{i}}\log p_{\theta}(x|z)=(x-\mu(z))\Sigma^{-1}\partial_{z_{i}}\mu(z).$ (25)
@@
$g_{ij}=\mathbb{E}[(\partial_{z_{i}}\log p(x|z)(\partial_{z_{j}}\log p(x|z))]$ $=\partial_{z_{i}}\mu^{T}(z)\Sigma^{-1}\underbrace{\mathbb{E}[(x-\mu)(x-\mu)^{T}]}_{=\Sigma}\Sigma^{-1}\partial_{z_{i}}\mu(z)$ (26)
@@ 5 Our work: complex decoder geometry
is the Hermitian metric for Riemannian metric $g$.
@@ 5.1 The Fisher information metric in the complex case
Generally it is the case that a Riemannian metric in a generative model has correspondende to a Fisher information metric. We argue that so does a Hermitian metric under the relation
@@
Remark. The Fisher information metric for a Euclidean Gaussian $X\sim\mathcal{N}(x;,\mu(x),\Sigma(x))$ is highly well known casella2002statistical, but not studied as often in the complex case. The primary difference in the complex case is the fact of $1/2$ appearing in the real-valued case, whereas here it does not due to the complex multivariate Gaussian density definition. It is possible to simplify the first term by taking
$(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)=\partial_{\alpha}\partial_{\overline{\beta}}(\mu^{\dagger}\Sigma^{-1}\mu).$ (37)
There is no immediate way to simplify $\partial_{\gamma}\Sigma$.
@@
$\ \ \ \ \ \ +v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\mu)(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma^{-1}v+v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\mu)v^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+v\Sigma^{-1}(\partial_{\alpha}\mu)v^{\dagger}Bv-v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\mu)t_{\overline{\beta}}$ (52)
@@ 5.2 The Fisher information metric and the Hessian of the KL divergence as a Kähler potential
$\partial\overline{\partial}\ \text{KL}\big(p(x|z)\parallel p(x|z^{\prime})\big)\Bigg|_{z^{\prime}=z}=2\text{Re}(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+\text{Tr}(\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}(\partial_{\overline{\beta}}\Sigma)),$ (67)
@@
5.3 A log-likelihood sum potential induces a metric close to the Fisher information metric in the continuum limit
@@
The primary goal of this section is that we will show there exists a Kähler potential that induces a Hermitian metric similarly to the true Fisher information metric. In particular, our method is more justified theoretically rather than simply ”replacing” terms in 36, since if we swap terms for things more computable, it will not necessarily hold that Kähler geometry exists. Thus, this section will attempt to justify we can approximate the Hermitian metric of 36 with something that is valid as a metric that is simultaneously more computable under assumptions.
We provide some motivation for this section. We begin by sampling a collection in the latent space. We refer to this as the previous samples. In this section, we anticipate a sampled point $z$ is sufficiently ”close” to a previously sampled mean, i.e. the previous sample is sufficiently dense. Under this assumption, we operate so that covariance is estimated sufficiently well under this density. Moreover, the metric we provide will be an expected value analog in the previous section, more or less. We will argue, under this density of points and due to exponential convergence properties, the weights of our expected value will be arranged in such a way that the weight of the ”close” points to the newly sampled point is significantly larger than those of previously sampled points far away. Thus our expectation will coincide with the terms in the previous section. Moreover, with all of this in mind, we still maintain validity of the metric. Thus, we have sound justification for approximately the metric of the previous section while maintaining efficiency.
@@
Now, using the exponential derivative and the product rule,
@@
where $\tilde{h}$ is a true target metric, which is an asymptotic identity, thus our methods serves as approximations. We provide empirical evidence of the above through Figure 4. We refer to Appendix B for full details.
@@
Figure 2: We show distributional equivalence by plotting the cumulative distribution of the $(1,1)$ real element of the first term Fisher information metric $(\partial_{\alpha}\mu)^{\dagger}\Sigma_{j}^{-1}(\partial_{\overline{\beta}}\mu)+(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma_{j}^{-1}(\partial_{\alpha}\mu)$ along the maximum index versus computing the expectation of $\mathbb{E}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi]$ using the softmax. The purpose of this plot is to convey the asymptotic equivalence of Lemmma 1, i.e. the softmax will be dominated by a singular weight.
@@
Figure 4: We present a cosine similarity between $\operatorname*{arg\,min}_{i}\Psi_{i}$ and $\operatorname*{arg\,max}_{i}\Psi_{i}$ on 50 posterior samples corresponding to $2\times\text{latent dim}$ indices (so one real, one imaginary) of our Kähler metric using pure weights of $w=1$ to replace the softmax function as of 93. The purpose of this plot is to convey the metric elements among these indices are proportional, thus the iterated expectation of 94 holds reasonably well. This figure was created using our MNIST experiment.
@@
Figure 5: We illustrate the condition as in Theorem 3, $\partial v\approx M\Sigma^{-1}v$, is reasonable through a cosine similarity. In particular, a similarity of $1\in[-1,+1]$ is most desirable. Ours is 0.793, which means there is reasonable similarity.
@@ 5.4 Inference
$\omega_{p}^{n}=d!\frac{i^{d}}{2^{d}}\zeta^{1}\wedge\overline{\zeta}^{1}\wedge\ldots\wedge\zeta^{d}\wedge\overline{\zeta}^{d},\ \ \ \ \ \zeta^{i}\wedge\overline{\zeta}^{i}=-2i\chi^{i}\wedge v^{i}$ (113)
we get $\omega_{p}^{d}=d!\text{Vol}_{g}$, but since $\omega^{n}=d!\text{det}(h)(i/2)^{d}dz^{1}\wedge d\overline{z}^{1}\wedge\ldots$, we have the result.
@@
$h(z)\in\Bigg\{h\ \ \Bigg|\ \ 1\approx\frac{\text{det}(h)}{\text{det}(\mathbb{E}_{w}[\partial\overline{\partial}\Psi]+\frac{1}{\rho^{2}}\text{Cov}(\partial\Psi,\overline{\partial}\Psi)))}\propto\text{det}(\partial\overline{\partial}\ \text{KL}(p(x|z)\parallel p(x|z^{\prime}))\Big|_{z^{\prime}=z})\Bigg\}$ (114)
where the determinant is real since $h$ is a valid Hermitian metric, and so
$h\in\Gamma(\text{Herm}^{+}(T^{1,0}M))\cap\mathcal{A}_{\delta},\ \ \ \mathcal{A}_{\delta}=\{h\ |\ \text{det}H_{\text{KL}}(z)-\text{det}H_{\Psi}(z)|<\delta\}.$ (115)
More specifically, we can facilitate the metric computation via faster approximations. For example, oftentimes, we can sample the diagonal only for computational benefit, and so
@@
The sampling is done at inference, so our metric surrogate is used in both training and inference. This loss encourages nondegeneracy ($h\approx 0$ will blow up), and extreme curvature ($h$ large will blow up negatively, but $z^{\dagger}z$ counteracts this). This loss helps mitigate ill-conditioning with the decoder metric. This loss has connections to existing literature but is largely new. nazari2023geometricautoencodersdecode proposes penalizing the induced metric determinant, but rather its variance in a non-variational setting, and so no sampling is done in accordance with the penalty. lee2022regularized presents isometric representation learning with a regularization penalty. It is nontrivial to compute the metric $h$ as part of the training procedure. Our solution to this is compute it every $K$ iterations and to attach the fixed, stored gradient to the gradient descent over iteration.
@@
   Sample batch, draw $\varepsilon\sim\mathcal{CN}(0,I)$;
@@
using a gradient-enhanced approach as of Yu_2022, which is a classical technique in physics-informed neural networks (PINNs).
Table 1: We illustrate outlier nearest neighbor distance (in the pixel space) scores using a baseline complex VAE. on our geometric Kähler training and sampling. All sampling is done with the aggregate posterior. Hyperparameter $\beta$ to scale the KL divergence is held constant here among ours and the baseline. Lower is better. 9xth denotes percentile, and $\%$ refers to what percent of data is outliers. The MNIST experiment is without metric normalization, and the CIFAR experiment is with normalization (subtract accumulated mean).
@@ 6 Conclusions
We have developed two Kähler potentials whose complex Hessian $\partial\overline{\partial}$ is the decoded Fisher information metric, one exact and one approximate. Our approximate metric remains a valid Kähler potential $K$, and it is efficient because the computational burden is displaced from automatic differentiation to an expectation. This expectation is also more efficient than using a nearest neighbor algorithm. The efficiency matters because constructing a metric of the latent space and sampling from this metric is nontrivial. Our work is primarily interesting because the intersection of complex geometry and ML is a largely unexplored area.
@@ 7 Acknowledgments
Appendix A Our Kähler potential and the Fisher metric analytic equivalence
@@
Let us examine the regime under the circumstance $\min_{i}\Psi_{i}$ is chosen. Let us assume parameter closeness and $\gamma I\preceq\Sigma_{i}$. With a slight lack of rigor, consider the following

### v4

@@ 2 Notation and conventions
to denote the complex Hessian with Dolbeault operators, i.e. it is a differential operator on smooth sections. We will moreover denote
@@
are the Wirtinger derivatives and $\mu$ is a mean vector (we are not saying all mixed Hessians vanish). We will show this is empirically supported in Appendix C, Figures 5, 6. In particular, we find our architectures work well with this criterion. We specifically do not restrict our decoder to be holomorphic since it is well known a real-valued function that is holomorphic over a connected and open domain is constant due to the Cauchy-Riemann equations. In practice, we will compute Wirtinger derivatives using real-valued surrogates in our VAEs for simplicity reasons. For complex integration, we will use the fact that complex integration has an equivalent form over $\mathbb{R}^{p}$
@@ 3 The complex variational autoencoder (VAE)
$\psi_{\text{Re}}=\frac{\sigma+\delta}{2\sigma+2\text{Re}(\delta)},\psi_{\text{Im}}=i\frac{\sqrt{\sigma^{2}-|\delta|^{2}}}{2\sigma+2\text{Re}(\delta)}.$ (11)
@@ 4 Euclidean VAEs unveil Riemannian structure
(under classical regularity conditions such as absolute continuity w.r.t. alternative measures, and slight abuse of notation on the subscripts). Thus the sampling is done with the constructed $G$ (or a valid surrogate or replacement). By sampling according to this distribution, they achieved a way of sampling with correspondence to the model geometry, which has shown empirical success in achieving high-fidelity generation chadebec2022geometricperspectivevariationalautoencoders. The arguments we will present will be in the context of Fisher information statistical theory. Our experiments, established in an alternative view, will not attempt to draw from a similar distribution for FID reasons as they do in this work.
@@
$\partial_{z_{i}}\log p_{\theta}(x|z)\propto(x-\mu(z))^{T}\Sigma^{-1}\partial_{z_{i}}\mu(z).$ (25)
@@
$g_{ij}=\mathbb{E}[(\partial_{z_{i}}\log p(x|z)(\partial_{z_{j}}\log p(x|z))]$ $=\partial_{z_{i}}\mu^{T}(z)\Sigma^{-1}\underbrace{\mathbb{E}[(x-\mu)(x-\mu)^{T}]}_{=\Sigma}\Sigma^{-1}\partial_{z_{j}}\mu(z)$ (26)
@@ 5 Our work: complex decoder geometry
is the Hermitian metric for Riemannian metric $g$. Note that there are various conventions, so we choose the above.
@@ 5.1 The Fisher information metric in the complex case
Generally it is the case that a Riemannian metric in a generative model has correspondence to a Fisher information metric. We argue that so does a Hermitian metric under the relation
@@
Remark. The Fisher information metric for a Euclidean Gaussian $X\sim\mathcal{N}(x;,\mu(x),\Sigma(x))$ is highly well known casella2002statistical, but not studied as often in the complex case. A key difference in the complex case is the fact of $1/2$ appearing in the real-valued case, whereas here it does not due to the complex multivariate Gaussian density definition.
@@
$\ \ \ \ \ \ +v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\mu)(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma^{-1}v+v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\mu)v^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\mu)v^{\dagger}Bv-v^{\dagger}\Sigma^{-1}(\partial_{\alpha}\mu)t_{\overline{\beta}}$ (51)
@@ 5.2 The Fisher information metric and the Hessian of the KL divergence as a Kähler potential
$\partial_{\alpha}\partial_{\overline{\beta}}\ \text{KL}\big(p(x|z)\parallel p(x|z^{\prime})\big)\Bigg|_{z^{\prime}=z}=2\text{Re}(\partial_{\alpha}\mu)^{\dagger}\Sigma^{-1}(\partial_{\overline{\beta}}\mu)+\text{Tr}(\Sigma^{-1}(\partial_{\alpha}\Sigma)\Sigma^{-1}(\partial_{\overline{\beta}}\Sigma)),$ (66)
@@
5.3 A log-likelihood sum potential induces a metric dependent on the Fisher information metric in the continuum limit
@@
The primary goal of this section is that we will show there exists a Kähler potential that induces a Hermitian metric that depends on the true Fisher information metric. In particular, our method is more justified theoretically rather than simply ”replacing” terms in 36, since if we swap terms for things more computable, it will not necessarily hold that Kähler geometry exists. Thus, this section will attempt to justify we can replace the Hermitian metric of 36 with something that is valid as a metric that is simultaneously more computable under assumptions.
We provide some motivation for this section. We begin by sampling a collection in the latent space. We refer to this as the previous samples. In this section, we anticipate a sampled point $z$ is sufficiently ”close” to a previously sampled mean, i.e. the previous sample is sufficiently dense. Under this assumption, we operate so that covariance is estimated sufficiently well under this density. Moreover, the metric we provide will be an expected value analog in the previous section, more or less. We will argue, under this density of points and due to exponential convergence properties, the weights of our expected value will be arranged in such a way that the weight of the ”distant” points to the newly sampled point is significantly larger than those of previously sampled points far away. Thus our expectation will coincide with the derivation in Appendix A, and has dependence on the Fisher metric. Moreover, we still maintain validity of the metric.
We provide some additional motivation. We will use Gaussian-like functions, but with positive quadratic form instead of the prototypical negative. Thus, distant points will be rewarded with high function value, whereas the global maximum of a Gaussian (or Gaussian-like) is (typically) at its mean. This is atypical so we motivate this for the following reasons. Our metric is valid under this positive quadratic form, whereas adding negative signs to the Kähler potential is not necessarily so. Moreover, our methods here match the methods as in Appendix A, thus the laws of total expectation and covariance hold between the two derivations. Were were to add negative signs to our Kähler potential, this is not the case. Thus, our methods serve more as replacements/proxies to the Fisher information metric rather than approximations. The edge posterior points will be rewarded in sampling among posterior variational sampling. In practice, we find this complements the loss of 118 well, and our computational results are nontrivially meaningful as in Table 1.
In our experiments, we will add to the loss
$\mathcal{L}=\mathbb{E}[z^{\dagger}z-\log\text{det}(h)].$ (68)
This will encourage a larger metric as $z$ is distant from the origin. This contributes to justification of our methods.
@@
Now, using the exponential (Wirtinger) derivative and the product rule,
@@
where $\tilde{h}$ is a true target metric, which is an asymptotic identity, thus our methods serves as approximations. We refer to Appendix B for full details.
@@
Figure 2: We show distributional equivalence by plotting the cumulative distribution of the $(1,1)$ real element of the first term Fisher information metric $(\partial_{\alpha}\mu)^{\dagger}\Sigma_{j}^{-1}(\partial_{\overline{\beta}}\mu)+(\partial_{\overline{\beta}}\mu)^{\dagger}\Sigma_{j}^{-1}(\partial_{\alpha}\mu)$ along the maximum index versus computing the expectation of $\mathbb{E}[\partial_{\alpha}\partial_{\overline{\beta}}\Psi]$ using the softmax. The purpose of this plot is to convey the asymptotic equivalence of Lemma 1, i.e. the softmax will be dominated by a singular weight.
@@
Figure 4: We illustrate the condition as in Theorem 3, $\partial v\approx M\Sigma^{-1}v$, is reasonable through a cosine similarity. In particular, a similarity of $1\in[-1,+1]$ is most desirable. Ours is 0.793.
@@ 5.4 Inference
$\omega_{p}^{d}=d!\frac{i^{d}}{2^{d}}\zeta^{1}\wedge\overline{\zeta}^{1}\wedge\ldots\wedge\zeta^{d}\wedge\overline{\zeta}^{d},\ \ \ \ \ \zeta^{i}\wedge\overline{\zeta}^{i}=-2i\chi^{i}\wedge v^{i},$ (113)
and we get $\omega_{p}^{d}=d!\text{Vol}_{g}$, but since $\omega^{d}=d!\text{det}(h)(i/2)^{d}dz^{1}\wedge d\overline{z}^{1}\wedge\ldots$, we have the result.
@@
$h(z)\in\Bigg\{h\ \ \Bigg|\ \ 1\approx\frac{\text{det}(h)}{\text{det}(\partial\overline{\partial}\ \text{KL}(p(x|z)\parallel p(x|z^{\prime}))\Big|_{z^{\prime}=z})}\sim\text{det}(\mathbb{E}[\mathbb{E}[\partial\overline{\partial}\Psi]+\frac{1}{\rho^{2}}\sum\text{Cov}(\partial\Psi,\overline{\partial}\Psi)])\Bigg\}.$ (114)
We have used $\sim$ to denote dependence, not necessarily proportionality. The determinant is real since $h$ is a valid Hermitian metric, and so
$h\in\Gamma(\text{Herm}^{+}(T^{1,0}M))\cap\mathcal{A}_{\delta},\ \ \ \mathcal{A}_{\delta}=\{\ h\ |\ \text{det}\ H_{\Psi}(z)>\mathbb{E}_{z\sim\text{uniform}}[h(z)],\mathbb{E}_{z\sim\text{uniform}}[H_{\text{KL}}(z)]\ \},$ (115)
where the inequality holds with high probability. More specifically, we can facilitate the metric computation via faster approximations. For example, oftentimes, we can sample the diagonal only for computational benefit, and so
@@
The sampling is done at inference, so our metric surrogate is used in both training and inference. We make a few remarks. This loss encourages nondegeneracy ($h\approx 0$ will blow up), and extreme curvature ($h$ large will blow up negatively, but $z^{\dagger}z$ counteracts this). This loss helps mitigate ill-conditioning with the decoder metric. Moreover, we previously discussed advantages of this loss in section 5.3. In particular, our proposed softmax-type expectation rewards distance, and so does this loss. In particular, this loss function especially supports our methods.
nazari2023geometricautoencodersdecode proposes penalizing the induced metric determinant, but rather its variance in a non-variational setting, and so no sampling is done in accordance with the penalty. lee2022regularized presents isometric representation learning with a regularization penalty. It is nontrivial to compute the metric $h$ as part of the training procedure. Our solution to this is compute it every $K$ iterations and to attach the fixed, stored gradient to the gradient descent over iteration.
@@
   Sample batch, draw $\varepsilon\sim\frac{1}{\sqrt{2}}\mathcal{N}(0,I)+\frac{i}{\sqrt{2}}\mathcal{N}(0,I)$ ;
@@
(note the cyclic property of the trace) using a gradient-enhanced approach as of Yu_2022, which is a classical technique in physics-informed neural networks (PINNs).
Table 1: We illustrate outlier nearest neighbor distance (in the feature space) scores using a baseline complex VAE. on our geometric Kähler training and sampling. All sampling is done with the aggregate posterior. Hyperparameter $\beta$ to scale the KL divergence is held constant here among ours and the baseline. Lower is better. 9xth denotes percentile, and $\%$ refers to what percent of data is outliers. The MNIST experiment is without metric normalization, and the CIFAR experiment is with normalization (subtract accumulated mean).
@@ 6 Conclusions
We have developed two Kähler potentials whose complex Hessian $\partial\overline{\partial}$ depends on the decoded Fisher information metric, one exact and one a proxy. Our approximate metric remains a valid Kähler potential $K$, and it is efficient because the computational burden is displaced from automatic differentiation to an expectation. This expectation is also more efficient than using a nearest neighbor algorithm. The efficiency matters because constructing a metric of the latent space and sampling from this metric is nontrivial. Our work is primarily interesting because the intersection of complex geometry and ML is a largely unexplored area.
@@ 7 Acknowledgments
Appendix A Connections between our Kähler potential and the Fisher metric with analytic equivalence
@@
Let us examine the regime under the circumstance of sufficient closeness. Let us assume parameter closeness and $\gamma I\preceq\Sigma_{i}$. With a slight lack of rigor, consider the following
@@ Appendix C Additional figures
Figure 7: We present a cosine similarity between $\operatorname*{arg\,min}_{i}\Psi_{i}$ and $\operatorname*{arg\,max}_{i}\Psi_{i}$ on 50 posterior samples corresponding to $2\times\text{latent dim}$ indices (so one real, one imaginary) of our Kähler metric using pure weights of $w=1$ to replace the softmax function as of 93. This figure is created without normalization via subtracting the mean, thus they are more easily correlated. This figure was created using our MNIST experiment.
Figure 8: We present a cosine similarity between $\operatorname*{arg\,min}_{i}\Psi_{i}$ and $\operatorname*{arg\,max}_{i}\Psi_{i}$ on 50 posterior samples corresponding to $2\times\text{latent dim}$ indices (so one real, one imaginary) of our Kähler metric using pure weights of $w=1$ to replace the softmax function as of 93. This figure is created with normalization via subtracting the mean, thus we capture correlation due to metric differences. This figure does not imply sampling according to our metric is unreasonable due to a lack of similarity. Indeed, this figure demonstrates a lack of similarity in regard to proportional scaling (which is anticipated), but recall our loss function encourages a high metric along the distant regions.